# Safe On-Orbit Dislodging of Deployable Structures via Robust Adaptive MPC

### v1

@@ Abstract
This paper proposes a novel robust adaptive model predictive controller for on-orbit dislodging. We consider the scenario where a servicer, equipped with a robot arm, must dislodge a client, a time-varying system composed of an underpowered jammed solar panel with a hybrid hinge system on a space station. Our approach leverages online set-membership identification to reduce the uncertainty to provide robust safety guarantees during dislodging despite bounded disturbances while balancing exploration and exploitation effectively in the parameter space. The feasibility of the developed robust adaptive MPC method is also examined through dislodging simulations and hardware experiments in zero-gravity and gravity environments, respectively. In addition, the advantages of our method are shown through comparison experiments with several state-of-the-art control schemes for both accuracy of parameter estimation and control performance.
@@ I Introduction
On-orbit failures have frequently occurred ever since the first artificial satellite was launched into space [1, 2, 3, 4] due to various unpredictable reasons, e.g., debris collision, temperature, electromagnetic, etc. These failures have caused countless economic and scientific losses, which have become an enormous challenge in the past decades. On-orbit servicing leveraging autonomous robotics systems is an increasingly popular research topic for rescuing on-orbit failure [5, 6, 7, 8]. Additional complexity arises from the parameter uncertainty of the client satellite, the unpredictable nature of component failures and the constraints posed by limited sensory feedback [9], actuation power [10], and real-time decision making [11].
@@
Dislodging an underpowered jammed component is one of the most common on-orbit maintenance tasks [12, 13, 14]. It involves safely manipulating objects that may be stuck or restrained due to various factors, such as microgravity [15], cold welding [16], stiction [13], or mechanical failure [2]. Dislodging a jammed component often requires applying significant, sudden force, which could potentially damage a delicate component. Few studies have considered using on-orbit robotics systems to address the problem. Previous work in [17] presented an adaptive control algorithm using a multi-robot system in a dislodging task for a solar panel. However, the algorithm did not consider safety constraints, which can lead to damage of both the servicer and client spacecraft. The risk of damage is exacerbated during the transient period when the adaptive controller is learning the system parameters. Although the adaptive controller is guaranteed to eventually learn the correct parameter, during the learning process, the adaptive controller can exert excessive force to dislodge the jammed component, potentially causing damage.
@@
Online parameter estimation can be leveraged to address the challenge of uncertainty during servicing missions to guarantee safety. Adaptive control for updating controller parameters using real-time measurement data has been extensively studied and gained popularity in recent decades [18, 19, 20]. Traditional adaptive control methods, such as Model Reference Adaptive Control (MRAC) [21], adjust controller parameters to ensure the system follows a desired reference model but have the drawback of being sensitive to unmodeled dynamics and typically cannot handle state and input constraints. Failure to enforce constraints on both the manipulator and manipulated object may lead to component damage during the servicing [22], which can cause significant costs and lost time due to the need to replace aerospace components via relaunch.
@@
Model predictive control (MPC) is a widely applied framework for managing uncertainties while ensuring constraint satisfaction and stability. In [23], an adaptive MPC algorithm for linear systems was introduced using a recursive approach, and its extension in [24] employed a tube-based framework to handle parametric uncertainty and additive disturbances in nonlinear systems. However, these methods rely on passive identification, preventing full exploitation of concurrent adaptation and control. A learning-based MPC in [25] used set-membership identification composed of two online phases: an adaptation phase with rigid tube-based robust MPC and a learning phase for uncertainty estimation, but only considers constant parameters, making it unsuitable for time-varying applications such as those found in aerospace. To address these limitations, this paper considers an adaptive controller that integrates set-membership identification to iteratively refine parameter bounds and employs robust MPC to ensure constraint enforcement and stability under worst-case conditions, ultimately improving performance as uncertainty estimates become less conservative.
In this paper, we propose a novel robust adaptive MPC with set-membership parameter estimation for the problem of safe dislodging. We demonstrate its implementation on our servicer during a dislodging task based on our client. The main contributions are summarized:
  • We transfer a specific hinge model with Euler-Lagrange form into a state-space equation and use our improved MPC algorithm for a dislodging task compared with two popular baselined methods to verify the feasibility and performance of our method in real applications.
Our robust adaptive MPC leverages dual-mode control [26] by incorporating time-varying parameter estimates into the MPC cost, balancing exploration which increases persistency of excitation to reduce uncertainty and exploitation which enhances control performance. Comparisons with a PID controller, adaptive control [17], and a state-of-the-art AMPC [25] demonstrate the superior performance and parameter estimation for time-varying systems. Moreover, our dual-state tube ensures safe parameter estimation within the initial set, remaining robust against future uncertainties.
@@
[…]
@@ II On-Orbit Dislodging Problem
In this section, we define the dislodging problem for an underpowered jammed solar panel, preventing it from unfolding through its passive actuation system. Considering both unknown parameters and control input from the underpowered jammed solar panel and manipulator, respectively, with their constraints during dislodging, our objective is to keep safety not only for the control action but also for parameter estimation.
@@ II-A Client and Servicer Agents
Client The client is a resident space object (RSO) with a malfunctioning solar panel. We assume the actuation system of the solar panel on its hinge is malfunctioning and cannot unfold the solar panel to the desired position. The solar panel is also jammed in a random position.
@@
We assume the frame $\mathcal{H}$ on the base of the client is stationary relative to frame $\mathcal{W}$ placed in the Earth as shown in Fig. 1. Considering the client can decouple the dynamics from the solar panel automatically by its inner stabilized system. This assumption holds during the whole process of dislodging.
Furthermore, we consider the client as a time-varying system in which the external environment may affect its stiffness properties, e.g., temperature, macro-gravity, vacuum conditions, radiation, etc., [27]. Also, we assume the dislodging contact location and orientation are ambiguous; the robot arm may not always be perpendicular to the rod, which means the $\alpha$ is not always $90^{\circ}$ and should also be time-varying as shown in Fig. 2. Hence, estimating unknown time-varying parameters for both the jammed component and grasping information during the dislodging under constraints becomes the biggest challenge in this task. We will pose this problem by dislodging the hinge from its initial position $\theta_{0}$ to a desired position $\theta^{*}$ with the calculated force applied by the servicer.
@@ II-B Dislodging Dynamics
In this section, we model the dynamics of the client which is underpowered and to be dislodged by the servicer. The malfunctioning solar panel is modeled as an underpowered hybrid hinge [28] system whose dynamics can be modeled by the Euler-Lagrange equation in joint space as
[…]
where $\theta_{k}$ is the angle of the hinge which has mass inertia $M\in\mathbb{R}_{+}$ and Coriolis and centripetal torque $C$. The spring-loaded hinge has stiffness $\mathcal{K}_{k}$ and viscous friction coefficient $\zeta_{k}$. The term
@@
is the Coulomb friction of the hybrid hinge with material-dependent friction coefficient $\mu$; $r_{e}$ denote the effective radius of the rotor and ϵ0+italic-$\epsilon\rightarrow 0\in\mathbb{R}_{+}$ to avoid the denominator to be zero. $\tau_{m_{k}}\in\{\tau_{m}\in\mathbb{R}:|\tau_{m}|\leq\delta_{\tau}\}$ denotes the external disturbance by the manipulator which is bounded through a constant value $\delta_{\tau}\in\mathbb{R}_{+}$ that affects the torque of the hinge.
The term $\mathbf{u}_{k}^{\top}\cdot\mathbf{r}_{\alpha}\times d_{g}$ describes the torque placed on the hinge by the servicer’s robotic arm where $\mathbf{r}_{\alpha}=\left[0,-\cos\alpha,\sin\alpha,0,0,0\right]^{\top}\in\mathbb{R}^{6}$ denotes the rotation vector for the end-effector (EE) of the robot arm on servicer relative to the perpendicular to the client; $\mathbf{u}_{k}{=}\begin{bmatrix}\mathbf{f}_{k},\boldsymbol{\tau}_{k}\end{bmatrix}^{\top}{\in}\ \mathbb{R}^{6}$ in which $\mathbf{f}_{k}{=}\left[{f_{x}}_{k},{f_{y}}_{k},{f_{z}}_{k}\right]^{\top}{\in}\mathbb{R}^{3}$ and $\boldsymbol{\tau}_{k}{=}\left[{\tau_{x}}_{k},{\tau_{y}}_{k},{\tau_{z}}_{k}\right]^{\top}{\in}\ \mathbb{R}^{3}$ denote the force and torque applied by the EE of the robot arm on the servicer to the dislodging contact point, respectively.
Figure 2: Diagram of the dislodging process that one servicer docks with a space station using its docking system and then dislodges a malfunctioning solar panel that is underpowered via its manipulator in space.
@@ II-C Parametric Uncertainty
where $\boldsymbol{\rho}_{k}^{*}\in\mathcal{P}$ and $\mathbf{H}_{\rho}\in\mathbb{R}^{n_{p}\times 3}$. Furthermore, we assume the change in parameters is bounded. Specifically, we assume there exists a value $\varphi_{\rho}\in\mathbb{R}_{+}$ that satisfies $\boldsymbol{\rho}_{k}^{*}\in\mathcal{P}$ for $k\in\mathbb{N}$ that
@@ II-D Safe Dislodging Constraints
where $\mathbf{x}_{l}=[\theta_{l},\dot{\theta}_{l}]^{\top}\in\mathbb{R}^{2\times 1}$ and $\mathbf{x}_{u}=[\theta_{u},\dot{\theta}_{u}]^{\top}\in\mathbb{R}^{2\times 1}$ denote the upper bound and lower bound of the state vector for all time $k\in\mathbb{R}_{+}$, respectively. The bounds $\mathbf{x}_{l}$ and $\mathbf{x}_{u}$ prevent overextending the hinge or causing the solar panel from colliding with another part of the client or servicer. Enforcing this constraint is challenging since it may require excessive force to stop the client before it overextends. Thus, by avoiding over-extension damage, we may cause excessive force damage.

### v2

@@ Abstract
This paper proposes a novel robust adaptive model predictive controller for on-orbit dislodging. We study orbit dislodging where a servicing spacecraft uses a robotic arm to free a jammed and unactuated solar panel mounted on a hybrid hinge that acts as a time-varying client on a space station. Our method couples online set-membership identification with a robust adaptive MPC to enforce safety under bounded disturbances. The controller explicitly balances exploration to excite the system and shrink uncertainty and exploitation to improve control performance through a dual-mode cost. The feasibility of the developed robust adaptive MPC method is also examined through dislodging simulations and hardware experiments in freefall and terrestrial laboratory environments, respectively. In addition, the advantages of our method are shown through comparison experiments with several state-of-the-art control schemes for both accuracy of parameter estimation and control performance.
NOMENCLATURE
$M$
Inertia matrix.
$C$
Coriolis and centripetal torque term.
$\mathcal{K}$
Stiffness parameter.
$\mathcal{D}$
Grasp factor related to contact geometry.
AA
State matrix of the linear parametric model.
$A_{i}$
Affine decomposition matrix of $A(\rho_{k})$.
$B$
Input matrix of the linear parametric model.
$B_{i}$
Affine decomposition matrix of $B(\rho_{k})$.
$\mathcal{P}$
Admissible parameter polytope.
$\mathbb{W}$
Bounded disturbance set.
$H_{\rho}$
Polytope matrices defining $\mathcal{P}$.
$H_{w}$
Polytope description of $\mathbb{W}$.
$\mathcal{U}$
Input constraint set.
$\mathcal{X}$
State constraint set.
$\mathbb{Z}$
Combined state–input constraint set.
$\tau_{f}$
Coulomb friction torque.
$\mathbf{K}$
Stabilizing feedback gain used in the tube policy.
$\mathbf{A}_{\text{cl}}$
Closed-loop map operator.
$\Gamma_{j,\varsigma|k}$
Nonnegative multipliers certifying tube constraints.
$\mathbb{X}_{T}$
Terminal set with robust positive invariance in tube variables.
$Q{,}R{,}\Upsilon$
Weighting matrix in the cost.
$\varOmega$
Collection of decision variables.
$\lambda$
Scaling factor used with $\mathbb{X}_{T}$ in the terminal condition.
$J_{N}$
Finite-horizon objective.
$\mu$
Friction coefficient.
$r_{e}$
Effective rotor radius.
$\epsilon$
Small positive smoothing constant.
$\tau_{m,k}$
External torque disturbance applied by the manipulator.
$\delta_{\tau}$
Bound on the external torque disturbance.
$\varsigma$
Step index along the prediction horizon.
$\alpha_{k}$
Angle between gripper direction and rod axis at time $k$.
$d_{g}$
Dislodging contact position measured from the hinge axis.
$r_{\alpha}$
Wrench direction vector used in the torque transfer term.
$\boldsymbol{\rho}$
Parameter vector.
$\zeta_{k}$
Viscous damping parameter.
$\phi_{\rho}$
Bound on the rate of change of $\boldsymbol{\rho}_{k}$.
$v_{\varsigma|k}$
Auxiliary input in the tube policy.
$c_{\varsigma|k}$
Auxiliary input used in the convex reformulation.
$\ell$
Stage and terminal cost function.
@@ I Introduction
On-orbit failures have frequently occurred ever since the first artificial satellite was launched into space [emme1966part, tafazoli2009study, flores2014review, luu2021review] due to multiple factors including orbital debris [putzar2008vulnerability], thermal cycling and temperature extremes [putz2020mechanical], radiation-induced degradation and surface charging [buitrago2024statistical], vacuum-induced cold welding and stiction [merstallinger2021cold], and electromagnetic interference [de2022simulation]. These failures have caused countless economic and scientific losses, which have become an enormous challenge in the past decades. On-orbit servicing leveraging autonomous robotics systems is an increasingly popular research topic for rescuing on-orbit failure [davis2019orbit, oestreich2021orbit, virgili2019convex, parikh2025safe]. Additionally, servicing missions confront two fundamental challenges. Model uncertainty from partially known and time-varying client parameters and failure modes degrades prediction and planning, while operational limits including limited sensing [boning2006identification], tight actuation bounds [ma2016adaptive], and strict real-time requirements shrink the feasible set of safe control actions [beaumet2011feasibility].
@@
Dislodging jammed components, where off-nominal friction temporarily raises the torque demand above the joint’s nominal requirement, remains a common on-orbit maintenance task [laing2009space, rivera2021study, jammedSolar]. It involves safely manipulating objects that may be stuck or restrained due to various factors, such as microgravity [higuchi2002unexpected], stiction [rivera2021study], or mechanical failure [tafazoli2009study]. Dislodging a jammed component often requires applying significant, sudden force, which could potentially damage a delicate component. Few studies have considered using on-orbit robotics systems to address the problem. Previous work in [gao2023autonomous] presented an adaptive control algorithm using a multi-robot system in a dislodging task for a solar panel. However, the algorithm did not consider safety constraints, which can lead to damage of both the servicer and client spacecraft. The risk of damage is exacerbated during the transient period when the adaptive controller is learning the system parameters. Although the adaptive controller is guaranteed to eventually learn the correct parameter, during the learning process, the adaptive controller can exert excessive force to dislodge the jammed component, potentially causing damage.
@@
Online parameter estimation can be leveraged to address the challenge of uncertainty during servicing missions to guarantee safety. Adaptive control for updating controller parameters using real-time measurement data has been extensively studied and gained popularity in recent decades [xu1993adaptive, ulrich2010modified, yu2019robust]. Traditional Model Reference Adaptive Control (MRAC) [nguyen2018model] updates controller parameters to track a reference model but is sensitive to unmodeled dynamics and does not natively enforce state or input constraints , which can be unreliable in practical implementation due to unmodeled dynamics. Failure to enforce constraints on both the manipulator and manipulated object may lead to component damage during the servicing [enomoto2023delicate], which can cause significant costs and lost time due to the need to replace aerospace components via relaunch.
Model predictive control (MPC) is widely adopted in terrestrial industries for explicit handling of constraints and for stability assurance, while within spacecraft guidance and control it is an limited but worthy of development and expanding research area with selective applications based on its strong safety and reliability. In [lorenzen2017adaptive], an adaptive MPC algorithm for linear systems was introduced using a recursive approach, and its extension in [kohler2021robust] employed a tube-based framework to handle parametric uncertainty and additive disturbances in nonlinear systems. However, these methods rely on passive identification, preventing full exploitation of concurrent adaptation and control. A learning-based MPC in [aboudonia2024adaptive] used set-membership identification composed of two online phases: an adaptation phase with rigid tube-based robust MPC and a learning phase for uncertainty estimation, but only considers constant parameters, making it unsuitable for time-varying applications such as those found in aerospace. To address these limitations, this paper considers an adaptive controller that integrates set-membership identification to iteratively refine parameter bounds and employs robust MPC to ensure constraint enforcement and stability under worst-case conditions, ultimately improving performance as uncertainty estimates become less conservative.
In this paper, we propose a novel robust adaptive MPC with set-membership parameter estimation for the problem of safe dislodging. We demonstrate its implementation on the servicer during a dislodging task based on the client. The main contributions are summarized:
  • We derive a state-space representation of the hinge model and implement the proposed MPC for the dislodging task. Comparative studies against two baseline methods evaluate feasibility and performance.
The robust adaptive MPC in this work leverages dual-mode control [wesselowski2003dual] by incorporating time-varying parameter estimates into the MPC cost, balancing exploration which increases persistency of excitation to reduce uncertainty and exploitation which enhances control performance. Comparisons with a PID controller, adaptive control [gao2023autonomous], and a state-of-the-art AMPC [aboudonia2024adaptive] demonstrate the superior performance and parameter estimation for time-varying systems. Moreover, the dual-state tube ensures safe parameter estimation within the initial set, remaining robust against future uncertainties.
@@
The sets of real numbers and positive real numbers are denoted by $\mathbb{R}$ and $\mathbb{R}_{+}$, respectively. The sequence of integers from $n_{1}$ to $n_{2}$ is represented by $\{n_{1},n_{2}\}_{\mathbb{N}}=\{n\in\mathbb{N}\mid n_{1}\leq n\leq n_{2}\}$. For a vector $\mathbf{p}$, $\mathbf{p}{[i]}$ refers to its $i\text{-th}$ element. The $i\text{-th}$ row of a matrix $\mathbf{A}$ is denoted by $\mathbf{A}]_{i}$. The Minkowski sum of two sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by $\mathcal{A}\oplus\mathcal{B}$, and the pontryagin set difference of two sets can be denoted as $\mathcal{A}\ominus\mathcal{B}$. $\boldsymbol{\gamma}\in\mathbb{R}^{N}$ denotes a column vector with $N$ rows that each row contains a scalar value $\gamma$. The convex hull of the elements of a set S is represented by co{S}. We use $\mathbf{u}_{k}$ for the (real, measured) vector at time $k$ and $\mathbf{x}_{\varsigma|k}$ for the vector predicted $\varsigma$ steps ahead at time $k$. The estimated term is denoted with a ($\hat{\penalty 10000\ }$) on the top and the upper bound of a variable using (¯\ $\bar{}\$) on the top.
@@ II On-Orbit Dislodging Problem
In this section, we define the dislodging problem for an unactuated jammed solar panel, preventing it from unfolding through its passive actuation system. Considering both unknown parameters and control input from the unactuated jammed solar panel and manipulator, respectively, with their constraints during dislodging, the objective is to keep safety not only for the control action but also for parameter estimation.
@@ II-A Client and Servicer Agents
Client The client is a free-flying resident space object (RSO) with a jamming solar panel. We assume the actuation system of the solar panel on its hinge is jamming and cannot unfold the solar panel to the desired position. The hinge is stronger than the truss on the solar panel. The solar panel is also jammed in a random position.
@@
We assume the frame $\mathcal{L}$ is a a standard Local-Vertical Local-Horizontal(LVLH) frame placed on the center of mass(CoM) of the space station of the Client. $\mathcal{W}$ denotes the earth-fixed frame, which should be located at the center of the Earth, as shown in Fig. 1. Considering the client is a free-flying aerospace system that can adjust it motion to decouple the dynamics from the solar panel during dislodging automatically by its inner stabilized system. This assumption holds during the whole process of dislodging.
Furthermore, we consider the client as a time-varying system in which the external environment may affect its stiffness properties, e.g., temperature, macro-gravity, vacuum conditions, radiation, etc., [tribble2020space]. Also, we assume the dislodging contact location and orientation are ambiguous; the robot arm may not always be perpendicular to the rod, which means the $\alpha$ is not always $90^{\circ}$ and should also be time-varying as shown in Fig. 2. Note that the docking system is a rigid connector, as shown in Fig. 2, which can ensure the Servicer keeps relative static with the space station during dislogding mission, and there is no relative decoupling dynamics needed. Hence, estimating unknown time-varying parameters for both the jammed component and grasping information during the dislodging under constraints becomes the biggest challenge in this task. We will pose this problem by dislodging the hinge from its initial position $\theta_{0}$ to a desired position $\theta^{*}$ with the calculated force applied by the servicer. The CoM is located at the position with distance from the pivot axis of the revolute joint as $d_{s}$. The dislodging contact position relative to the revolute joint is $d_{g}$. The angle between the z-axis of the end-effector (EE) of the robot arm on Servicer and the central axis of the rod denotes $\alpha$, and the angle between the z-axis of EE and its perpendicualr axis relative to the central axis of the solar panel denotes $\phi$ in which $\phi=90^{\circ}-\alpha$ as shown in Fig. 2.
@@ II-B Dislodging Dynamics
In this section, we model the dynamics of the client which is unactuated and to be dislodged by the servicer. The functioning-well solar panel is modeled as an unactuated hybrid hinge [fang2012shape] system whose dynamics can be modeled by the Euler-Lagrange equation in joint space as
$\tau_{h}=M{{\ddot{\theta}}_{k}}+\Lambda\dot{\theta}_{k}+\mathcal{K}_{k}\theta_{k}+\tau_{f}(\dot{\theta}_{k})+\tau_{m_{k}},$ (1)
where $\tau_{h}=\mathbf{u}_{k}^{\top}\cdot\mathbf{r}_{\alpha}\times d_{g}\in\mathbb{R}_{+}$ denotes the torque applied on the hinge on the client by the servicer during dislodging, in which $\mathbf{r}_{\alpha}=\left[0,-\cos\alpha,\sin\alpha,0,0,0\right]^{\top}\in\mathbb{R}^{6}$ denotes the rotation vector for the EE of the robot arm on servicer relative to the perpendicular to the client; $\mathbf{u}_{k}{=}\begin{bmatrix}\mathbf{f}_{k},\boldsymbol{\tau}_{k}\end{bmatrix}^{\top}{\in}\ \mathbb{R}^{6}$ in which $\mathbf{f}_{k}{=}\left[{f_{x}}_{k},{f_{y}}_{k},{f_{z}}_{k}\right]^{\top}{\in}\mathbb{R}^{3}$ and $\boldsymbol{\tau}_{k}{=}\left[{\tau_{x}}_{k},{\tau_{y}}_{k},{\tau_{z}}_{k}\right]^{\top}{\in}\ \mathbb{R}^{3}$ denote the force and torque applied by the EE of the robot arm on the servicer to the dislodging contact point, respectively. $\theta_{k}$ denotes the angle of the hinge; $M\in\mathbb{R}_{+}$ denotes the mass inertia of the hinge. $\Lambda=C_{k}+\mathcal{\zeta}_{k}$ in which $C\in\mathbb{R}_{+}$ is the Coriolis and centripetal torque of the solar panel with friction torque limiter [^1] Structure of friction torque limiter, by MACH III CLUTCH. and $\mathcal{\zeta}_{k}\in\mathbb{R}_{+}$ is the viscous friction coefficient of the hybrid hinge.. The spring-loaded hinge has stiffness $\mathcal{K}_{k}$ and viscous friction coefficient $\zeta_{k}$. The term
@@
is the Coulomb friction of the hybrid hinge based on the dataset [^2] Supported material: Hinge Dataset. we collected with material-dependent friction coefficient $\mu$; $r_{e}$ denote the effective radius of the rotor and $\epsilon\rightarrow 0\in\mathbb{R}_{+}$ to avoid the denominator to be zero. $\tau_{m_{k}}\in\{\tau_{m}\in\mathbb{R}:|\tau_{m}|\leq\delta_{\tau}\}$ denotes the external disturbance by the manipulator which is bounded through a constant value $\delta_{\tau}\in\mathbb{R}_{+}$ that affects the torque of the hinge. Note that this hybrid hinge is a torsional design, where a coaxial helical spring produces a restoring torque modeled by a linear stiffness law.
Figure 2: Diagram of the dislodging process that one servicer docks with a space station using its docking system and then dislodges a jamming solar panel that is unactuated via its manipulator in space.
@@ II-C Parametric Uncertainty
where $\boldsymbol{\rho}_{k}^{*}\in\mathcal{P}$ and $\mathbf{H}_{\rho}\in\mathbb{R}^{n_{p}\times 3}$ in which $n_{p}$ denotes the number of the hyperplanes used to define the known bounded polytope $\mathcal{P}$. Furthermore, we assume the change in parameters is bounded. Specifically, we assume there exists a value $\varphi_{\rho}\in\mathbb{R}_{+}$ that satisfies $\boldsymbol{\rho}_{k}^{*}\in\mathcal{P}$ for $k\in\mathbb{N}$ that
@@
We thus clarify that the system state $\boldsymbol{\theta}_{k}$ describes only the hinge angle, while the off-pointing geometry of the Servicer’s EE $\alpha_{k}$ is treated as an unknown, time-varying parameter handled within the estimation vector $\boldsymbol{\rho}_{k}$.
@@ II-D Safe Dislodging Constraints
where $\mathbf{x}_{l}=[\theta_{l},\dot{\theta}_{l}]^{\top}\in\mathbb{R}^{2\times 1}$ and $\mathbf{x}_{u}=[\theta_{u},\dot{\theta}_{u}]^{\top}\in\mathbb{R}^{2\times 1}$ denote the upper bound and lower bound of the state vector for all time $k\in\mathbb{R}_{+}$, respectively. The bounds $\mathbf{x}_{l}$ and $\mathbf{x}_{u}$ prevent overextending the hinge or causing the solar panel from colliding with another part of the client or servicer. Enforcing this constraint is challenging, since preventing over-extension may require large interaction forces that could overstress the Client structure.